# Early Exploration of a Flexible Framework for Efficient Quantum Linear Solvers in Power Systems

### v2

@@ Abstract
This is a preprint version of the work. This work has been accepted by the IEEE. ©2024 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, including reprinting/republishing this material for advertising or promotional purposes, collecting new collected works for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ I Introduction
quantum advantage. This is a preprint version of the work. This work has been accepted by the IEEE. ©2024 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, including reprinting/republishing this material for advertising or promotional purposes, collecting new collected works for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@
Given all these challenges, the authors propose a flexible simulator-based framework designed for tackling power system problems. This framework leverages quantum simulators, which are dedicated to the evaluation of quantum circuits on classical quantum computers. The use of simulators can facilitate feasibility studies in this area, targeting more realistic problem settings and making it more accessible to a broader range of researchers. Once fault-tolerant quantum computers are available, researchers can then migrate their existing work
to the real hardware, further advancing the field. This is a preprint version of the work. This work has been accepted by the IEEE. ©2024 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, including reprinting/republishing this material for advertising or promotional purposes, collecting new collected works for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ II-C Embedded techniques in SV-Sim simulator
Another source of time inefficiency stems from latency arising from communication between multiple computational units during parallel computing. On one hand, SV-Sim employs the so-called “PGAS-based SHMEM” communication model,
which provides enhanced support for software governing direct links between GPUs compared to the traditional MPI routine. On the other hand, researchers have also developed a novel method to minimize unnecessary data migration [6]. This is a preprint version of the work. This work has been accepted by the IEEE. ©2024 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, including reprinting/republishing this material for advertising or promotional purposes, collecting new collected works for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ III-A Procedures for applying HHL on practical problems
Let us consider the linear system $\mathbf{A}\mathbf{x}=\mathbf{b}$ in an AC power flow problem. $\mathbf{A}$ is assumed to be Hermitian and an easy conversion can be conducted if it is not. HHL algorithm aims to find the normalized solution vector $|\mathbf{x}\rangle$ by $|\mathbf{x}\rangle=\mathbf{A}^{-1}|\mathbf{b}\rangle=\sum_{i}\lambda_{i}^{-1}b_{i}|\mathbf{v_{i}}\rangle$, where $\lambda_{i}$ and $\mathbf{v_{i}}$ are the $i^{th}$ eigenvalue and eigenvector of $\mathbf{A}$, respectively, and $|\mathbf{b}\rangle=\sum_{i}b_{i}|\mathbf{v_{i}}\rangle$ is the normalized vector $\mathbf{b}$ with decomposition. The decomposition of $\mathbf{b}$ in the eigenbasis of $\mathbf{A}$ is achieved with quantum phase estimation (QPE), as shown in the circuit in Fig. 5. The complete procedures used to solve the linear system from the power flow problems are following.
Pratical HHL procedures Goal: Providing $\mathbf{A}$ and $\mathbf{b}$ in a linear system $\mathbf{A}\mathbf{x}=\mathbf{b}$, use HHL to solve for solution vector $\mathbf{x}$. Process: 1. First, normalize $\mathbf{b}$ to get $|\mathbf{b}\rangle$ and record $\|\mathbf{b}\|$, then check the properties of matrix $\mathbf{A}$: (a) if the condition number is too high, use a preconditioner (b) if the dimension is not a power of 2, expand it to the nearest power of 2 (c) if the matrix is not Hermitian (e.g., due to the preconditioner), make it Hermitian: $\left[\begin{array}[]{cc}\mathbf{0}&\mathbf{A}\\ \mathbf{A}^{\dagger}&\mathbf{0}\end{array}\right]\left[\begin{array}[]{c}\mathbf{0}\\ |\mathbf{x}\rangle\end{array}\right]=\left[\begin{array}[]{c}|\mathbf{b}\rangle\\ \mathbf{0}\end{array}\right].$ 2. Generate the HHL circuit as in Figure 5. 3. Compute $|\mathbf{x}\rangle$ by measuring the final state $|\vec{\psi}\rangle$ from the circuit. If a simulator is used, the statevector of $|\psi\rangle$ can be read directly. if a real quantum computer is used, the statevector of $|\psi\rangle$ has to be reconstructed from state tomography. 4. Retrieve $\mathbf{x}$ from $|\mathbf{x}\rangle$ and return it. (a) Compute $\|\mathbf{A}^{-1}\left|\mathbf{b}\right\rangle\|=|\lambda_{0}|\sqrt{P(\text{\small meas. ancilla and get 1})}$ where $|\lambda_{0}|\leq|\lambda_{i}|$ for all $i\neq 0$, and $\lambda_{i}$ are eigenvalues of the matrix input in the circuit. (b) Finally, $\mathbf{x}=\|\mathbf{A}^{-1}\left|b\right\rangle\|\cdot\|\mathbf{b}\|\cdot|\mathbf{x}\rangle$.
@@ III-B Resources estimation and scalability
available GPU nodes. However, as shown in Table I, the current bottleneck is in HHL circuit generation. This is a preprint version of the work. This work has been accepted by the IEEE. ©2024 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, including reprinting/republishing this material for advertising or promotional purposes, collecting new collected works for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ IV Numerical experiments
Experiments were conducted on Perlmutter supercomputer using a single NVIDIA A100 GPU. We will use multiple GPUs to accommodate the increased computational demands for future large test cases. In our study, HHL circuits for both the 14-bus and 30-bus cases were executed on both Qiskit Aer and the SV-Sim simulator. The detailed results are presented in Table I. A highlight is the impact of the GS preconditioner on the 30-bus case. It can reduce the condition number of the 30-bus system from 492.5 to 109.3. This reduction had the effect of lowering the value of $n_{\text{qpe}}$ from 10 to 8, as elaborated in the entry marked “$*{}^{*}$” in the table.
@@ V Conclusion and future work
TABLE I: Experiment results on 14-bus case, 30-bus case, and 30-bus case with GS preconditioner ($*{}^{*}$)
14-bus 30-bus $*{}^{*}$

### v3

@@ Abstract
M. Zheng, Y. Chen, X. Yang and A. Li, ”Early Exploration of a Flexible Framework for Efficient Quantum Linear Solvers in Power Systems,” 2024 IEEE Power & Energy Society General Meeting (PESGM), Seattle, WA, USA, 2024, pp. 1-5, doi: 10.1109/PESGM51994.2024.10688916.
©2024 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, including reprinting/republishing this material for advertising or promotional purposes, collecting new collected works for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ I Introduction
quantum advantage. M. Zheng, Y. Chen, X. Yang and A. Li, ”Early Exploration of a Flexible Framework for Efficient Quantum Linear Solvers in Power Systems,” 2024 IEEE Power & Energy Society General Meeting (PESGM), Seattle, WA, USA, 2024, pp. 1-5, doi: 10.1109/PESGM51994.2024.10688916.
©2024 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, including reprinting/republishing this material for advertising or promotional purposes, collecting new collected works for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@
Given all these challenges, the authors propose a flexible simulator-based framework designed for tackling power system problems. This framework leverages quantum simulators, which are dedicated to the evaluation of quantum circuits on classical quantum computers. The use of simulators can facilitate feasibility studies in this area, targeting more realistic problem settings and making it more accessible to a broader range of researchers. Once fault-tolerant quantum computers are available, researchers can then migrate their existing work to the real hardware, further advancing the field.
M. Zheng, Y. Chen, X. Yang and A. Li, ”Early Exploration of a Flexible Framework for Efficient Quantum Linear Solvers in Power Systems,” 2024 IEEE Power & Energy Society General Meeting (PESGM), Seattle, WA, USA, 2024, pp. 1-5, doi: 10.1109/PESGM51994.2024.10688916.
©2024 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, including reprinting/republishing this material for advertising or promotional purposes, collecting new collected works for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ II-C Embedded techniques in SV-Sim simulator
Another source of time inefficiency stems from latency arising from communication between multiple computational units during parallel computing. On one hand, SV-Sim employs the so-called “PGAS-based SHMEM” communication model, which provides enhanced support for software governing direct links between GPUs compared to the traditional MPI routine.
On the other hand, researchers have also developed a novel method to minimize unnecessary data migration [6]. M. Zheng, Y. Chen, X. Yang and A. Li, ”Early Exploration of a Flexible Framework for Efficient Quantum Linear Solvers in Power Systems,” 2024 IEEE Power & Energy Society General Meeting (PESGM), Seattle, WA, USA, 2024, pp. 1-5, doi: 10.1109/PESGM51994.2024.10688916.
©2024 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, including reprinting/republishing this material for advertising or promotional purposes, collecting new collected works for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ III-A Procedures for applying HHL on practical problems
Let us consider the linear system $\mathbf{A}\mathbf{x}=\mathbf{b}$ in an AC power flow problem.
$\mathbf{A}$ is assumed to be Hermitian and an easy conversion can be conducted if it is not. HHL algorithm aims to find the normalized solution vector $|\mathbf{x}\rangle$ by $|\mathbf{x}\rangle=\mathbf{A}^{-1}|\mathbf{b}\rangle=\sum_{i}\lambda_{i}^{-1}b_{i}|\mathbf{v_{i}}\rangle$, where $\lambda_{i}$ and $\mathbf{v_{i}}$ are the $i^{th}$ eigenvalue and eigenvector of $\mathbf{A}$, respectively, and $|\mathbf{b}\rangle=\sum_{i}b_{i}|\mathbf{v_{i}}\rangle$ is the normalized vector $\mathbf{b}$ with decomposition. The decomposition of $\mathbf{b}$ in the eigenbasis of $\mathbf{A}$ is achieved with quantum phase estimation (QPE), as shown in the circuit in Fig. 5. The complete procedures used to solve the linear system from the power flow problems are following.
{mdframed}
[linecolor=black,linewidth=1pt] Practical HHL procedures
Goal: Providing $\mathbf{A}$ and $\mathbf{b}$ in a linear system $\mathbf{A}\mathbf{x}=\mathbf{b}$, use HHL to solve for solution vector $\mathbf{x}$.
Process:
  1. 1.
    First, normalize $\mathbf{b}$ to get $|\mathbf{b}\rangle$ and record $\|\mathbf{b}\|$, then check the properties of matrix $\mathbf{A}$:
    1. (a)
      if the condition number is too high, use a preconditioner
    2. (b)
      if the dimension is not a power of 2, expand it to the nearest power of 2
    3. (c)
      if the matrix is not Hermitian (e.g., due to the preconditioner), make it Hermitian:
      $\left[\begin{array}[]{cc}\mathbf{0}&\mathbf{A}\\ \mathbf{A}^{\dagger}&\mathbf{0}\end{array}\right]\left[\begin{array}[]{c}\mathbf{0}\\ |\mathbf{x}\rangle\end{array}\right]=\left[\begin{array}[]{c}|\mathbf{b}\rangle\\ \mathbf{0}\end{array}\right].$
  2. 2.
    Generate the HHL circuit as in Figure 5.
  3. 3.
    Compute $|\mathbf{x}\rangle$ by measuring the final state $|\mathbf{\psi}\rangle$ from the circuit.
    • If a simulator is used, the statevector of $|\mathbf{\psi}\rangle$ can be read directly.
    • if a real quantum computer is used, the statevector of $|\mathbf{\psi}\rangle$ has to be reconstructed from state tomography.
  4. 4.
    Retrieve $\mathbf{x}$ from $|\mathbf{x}\rangle$ and return it.
    1. (a)
      Compute
      $\|\mathbf{x}\|=|\lambda_{0}|\cdot\sqrt{P(\text{measure ancilla and get 1})},$
      where $|\lambda_{0}|\leq|\lambda_{i}|$ for all $i\neq 0$, and $\lambda_{i}$ are eigenvalues of the matrix input in the circuit.
    2. (b)
      Finally, $\mathbf{x}=\|\mathbf{x}\|\cdot\|\mathbf{b}\|\cdot|\mathbf{x}\rangle$.
@@ III-B Resources estimation and scalability
available GPU nodes. However, as shown in Table I, the current bottleneck is in HHL circuit generation. M. Zheng, Y. Chen, X. Yang and A. Li, ”Early Exploration of a Flexible Framework for Efficient Quantum Linear Solvers in Power Systems,” 2024 IEEE Power & Energy Society General Meeting (PESGM), Seattle, WA, USA, 2024, pp. 1-5, doi: 10.1109/PESGM51994.2024.10688916.
©2024 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, including reprinting/republishing this material for advertising or promotional purposes, collecting new collected works for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ IV Numerical experiments
Experiments were conducted on Perlmutter supercomputer using a single NVIDIA A100 GPU. We will use multiple GPUs to accommodate the increased computational demands for future large test cases. In our study, HHL circuits for both the 14-bus and 30-bus cases were executed on both Qiskit Aer and the SV-Sim simulator. The detailed results are presented in Table I. A highlight is the impact of the GS preconditioner on the 30-bus case. It can reduce the condition number of the 30-bus system from 492.5 to 109.3. This reduction had the effect of lowering the value of $n_{\text{qpe}}$ from 10 to 8, as elaborated in the entry marked “30-buses” in the table.
@@ V Conclusion and future work
TABLE I: Experiment results on 14-bus case, 30-bus case, and 30-bus case with GS preconditioner (30-bus)
14-bus 30-bus 30-bus